OpenStudy (anonymous):

Sixty percent of the populace were iconoclasts; the rest were traditionalists. If 16,000 were traditionalists, how many were iconoclasts ?

OpenStudy (anonymous):

Make an equation. 60% are iconoclasts. 100% - 60% = 40% are traditionalists. Total population is 16,000 + x. 0.6(16,000 + x) + 16,000 = 16000 + x Solve for x.

OpenStudy (anonymous):

Here's a cool way to do it.. Perhaps nonstandard, though 40% are traditionalists, right? so 40% is 16,000 divide by 4 to get 10% is 4,000 multiply by 6 to get 60% is 24,000....there's your answer. 24,000

OpenStudy (anonymous):

Steps: 0.6(16,000) + 0.6x = x (subtracted 16000 from both sides, also distributed the 0.6 0.6(16,000) = 0.4x (subtracted 0.6x from both sides) \[\frac{0.6(16,000)}{0.4} = x = 24,000\]

OpenStudy (anonymous):

Another way to do this is with a ratio: \[\frac{16,000}{0.4} = \frac{x}{0.6}\] solve for x, this is fast

OpenStudy (anonymous):

I think most students would figure out what number 16000 is 40% of by solving 16000 = .4x getting x = 40000 Then find 60% of 40000 as 0.6 * 40000= 24000

OpenStudy (anonymous):

Lots of different ways. LOL

OpenStudy (anonymous):

yeah lol well this guy is set